What is 4 1/3 divided by 11 4/5?
How do you solve this problem? Do you cross multiply...Please explain step by step

Answer by rfer(16322)

(4 1/3)/11 4/5)=
change to improper fractions
(13/3)/(59/5)=
invert and multiply
13/3*5/59=65/177